# Diverse beam search to find densest-known planar unit distance graphs

### v1

@@ 2.1. The Moser Lattice
$M_{L}=\{a\cdot 1+b\cdot\omega_{2}+c\cdot\omega_{3}+d\cdot\omega_{2}\omega_{3}\mid a,b,c,d\in\mathbb{Z}\}=\mathbb{Z}\langle 1,\omega_{2},\omega_{3},\omega_{2}\omega_{3}\rangle\leq(\mathbb{C},+),$
where $\omega_{2}=\frac{1+i\sqrt{3}}{2}$ and $\omega_{3}=\frac{5+i\sqrt{11}}{6}$.
@@
    $\{1,\omega_{2},\omega_{3},\omega_{2}\omega_{3}\}$ is a basis for $M_{L}$.
@@
Firstly, consider the minimal polynomials of $\omega_{2}$ and $\omega_{3}$ over $\mathbb{Q}$. As we have $\omega_{2}^{2}-\omega_{2}+1=0$ and $\omega_{3}^{2}-\frac{5}{3}\omega_{3}+1=0$, the two minimal polynomials are quadratic and coprime, which implies that $[\mathbb{Q}[\omega_{2},\omega_{3}]:\mathbb{Q}]=4$. That is, the degree of the extension is 4.
Secondly, observe that the 4-element set $\{1,\omega_{2},\omega_{3},\omega_{2}\omega_{3}\}$ generates the $\mathbb{Q}$-vector space $\mathbb{Q}[\omega_{2},\omega_{3}]$. Thus, this set is a basis for $M_{L}$.
Lastly, define the map $\varphi:\mathbb{Z}^{4}\rightarrow M_{L}$, where $\varphi((a,b,c,d))=a\cdot 1+b\cdot\omega_{2}+c\cdot\omega_{3}+d\cdot\omega_{2}\omega_{3}$. We show that this map is an isomorphism. The injectivity of $\varphi$ is clear from the linear independence of $1$, $\omega_{2}$, $\omega_{3}$, and $\omega_{2}\omega_{3}$. The surjectivity of $\varphi$ comes from the definition of $M_{L}$. Thus, $\mathbb{Z}^{4}\cong M_{L}$. ∎
@@
Define the map $\varphi:\mathbb{Z}^{4}\rightarrow M_{L}$, where $\varphi((a,b,c,d))=a\cdot 1+b\cdot\omega_{2}+c\cdot\omega_{3}+d\cdot\omega_{2}\omega_{3}$. For a UDG $U=(V,E)$ with vertices on $M_{L}$, the matrix form of $U$ is represented as $U=[\varphi^{-1}(v_{1}),\varphi^{-1}(v_{2}),\ldots,\varphi^{-1}(v_{n})]$, where the vertex set $V=\{v_{1},v_{2},\ldots,v_{n}\}$ and the vertices are ordered.
@@
Take an arbitrary element $z=a+b\omega_{2}+c\omega_{3}+d\omega_{2}\omega_{3}\in M_{L}$ with $a,b,c,d\in\mathbb{Z}$. It is of unit length if and only if
@@
$=(a+b\omega_{2}+c\omega_{3}+d\omega_{2}\omega_{3})\overline{(a+b\omega_{2}+c\omega_{3}+d\omega_{2}\omega_{3})}$
@@
$\text{with }\mathbf{v}:=\begin{pmatrix}a\\ b\\ c\\ d\end{pmatrix}\text{ and }A:=\begin{pmatrix}1&\frac{1}{2}&\frac{5}{6}&\frac{5}{12}\\ \frac{1}{2}&1&\frac{5}{12}&\frac{5}{6}\\ \frac{5}{6}&\frac{5}{12}&1&\frac{1}{2}\\ \frac{5}{12}&\frac{4}{5}&\frac{1}{2}&1\end{pmatrix}\text{ we have }p(a,b,c,d)=\mathbf{v}^{T}A\mathbf{v}.$
@@ 2.2. Canonization
Consider some matrix $[[p_{1}],[p_{2}],...,[p_{n}]]$ where each $[p_{i}]$ is a point on the Moser lattice, $[p_{i}]=[a_{i},b_{i},c_{i},d_{i}]=a_{i}+b_{i}\cdot\omega_{2}+c_{i}\cdot\omega_{3}+d_{i}\cdot\omega_{2}\omega_{3}$. Together, this array of points will constitute the vertex set of a planar unit distance graph. Immediately, there are $n!$ identical copies of this graph according to the permutations of this array. This may be addressed by treating this list as a set object.
@@
In the Moser lattice, we find $12$ representations of a graph through rotations by $\frac{\pi}{3}$ radians and reflections swapping the pair of generators $\omega_{2}$ and $\omega_{3}$ along with the pair $\omega_{2}\omega_{3}$ and $1$.
@@
$1=e^{0},\omega_{2}=e^{\frac{\pi i}{3}},\omega_{3}=e^{i\cdot 2\arcsin(\frac{1}{\sqrt{12}})}$
@@
$a\mapsto a\cdot\omega_{2},b\cdot\omega_{2}\mapsto-b+b\cdot\omega_{2},c\cdot\omega_{3}\mapsto c\cdot\omega_{2}\omega_{3},\text{ and }d\cdot\omega_{2}\omega_{3}\mapsto-d\cdot\omega_{3}+d\cdot\omega_{2}\omega_{3}.$
@@ 4.3.1. Operation 1: Single Edge Addition
Given a parent graph $G$ and a vertex $v\in V(G)$, generate $V(G^{\prime})=V(G)\cup v^{\prime}$ for a $v^{\prime}$ in the set of generators times ±$\pm 1$, {±$\{\pm 1,\pm\omega_{2},\pm\omega_{3},\pm\omega_{2}\omega_{3}\}$.
@@ 4.4. Parents
Finding clusters of edge-dense graphs using backtracking beam search relies also on finding subgraphs which lead to larger maximal graphs. These graphs can also therefore be found as parents of previously discovered graphs in the search. The algorithm for discovering parents is relatively simple and low-cost, and is as follows.
Given a graph $G$, we obtain the parent set of $G$ by creating some $V(G^{\prime})=V(G)-v$ for all $v\in V(G)$. The set of unique canonized parents of $G$ is thus at most the size of $G$, and some subgraphs may reduce to the same canonized graph.
@@ 6.2. The Moser Ring
$(a\cdot 1+b\cdot\omega_{2}+c\cdot\omega_{3}+d\cdot\omega_{2}\omega_{3})(a^{\prime}\cdot 1+b^{\prime}\cdot\omega_{2}+c^{\prime}\cdot\omega_{3}+d^{\prime}\cdot\omega_{2}\omega_{3}),$
for all $a,b,c,d,a^{\prime},b^{\prime},c^{\prime},d^{\prime}\in\mathbb{Z}$. $M_{R}$ is an extension ring $\mathbb{Z}[\omega_{2},\omega_{3}]$, where $\omega_{2}^{2}=\omega_{2}^{2}-1$, $66\omega_{3}^{2}=10\omega_{3}-$ 6, and $\omega_{3}^{2}=\frac{5}{3}\omega_{3}-1$.
@@
$B=\{1,\omega_{2},\omega_{3},\omega_{2}\omega_{3}\},$
where $\omega_{2}=\frac{1+i\sqrt{3}}{2}$ and $\omega_{3}=\frac{5+i\sqrt{11}}{6}$. Similarly to the matrix form of UDGs with vertices in the Moser lattice, as given in Definition 2.3, every element $z\in M_{R}$ of the Moser ring can be written as $z=\frac{\sum_{b_{i}\in B}a_{i}b_{i}}{3^{k}}$ for some $(a_{0},a_{1},a_{2},a_{3},k)\in\mathbb{Z}^{5}$. Moreover, simplifying fractions, we get a unique such representation. Thus, for a UDG with $n$ vertices in the Moser ring, we could represent it by an $n\times 5$ integer matrix. All the beam search algorithms would be similar to the Moser lattice, except for the get-children function. We have a multiplication operator besides an addition operator for creating children of a UDG. Therefore, we potentially can get more children in the Moser ring than in the Moser lattice.

### v2

@@ 2.1. The Moser Lattice
Notation 2.1.
For a positive integer $t\in\mathbb{Z}_{>0}$, we let
$\omega_{t}=\exp(i\cdot\arccos(1-1/2t)).$
@@
$M_{L}=\{a\cdot 1+b\cdot\omega_{1}+c\cdot\omega_{3}+d\cdot\omega_{1}\omega_{3}\mid a,b,c,d\in\mathbb{Z}\}=\mathbb{Z}\langle 1,\omega_{1},\omega_{3},\omega_{1}\omega_{3}\rangle\leq(\mathbb{C},+),$
@@
    $\{1,\omega_{1},\omega_{3},\omega_{1}\omega_{3}\}$ is a basis for $M_{L}$.
@@
Firstly, consider the minimal polynomials of $\omega_{1}$ and $\omega_{3}$ over $\mathbb{Q}$. As we have $\omega_{1}^{2}-\omega_{1}+1=0$ and $\omega_{3}^{2}-\frac{5}{3}\omega_{3}+1=0$, the two minimal polynomials are quadratic and coprime, which implies that $[\mathbb{Q}[\omega_{1},\omega_{3}]:\mathbb{Q}]=4$. That is, the degree of the extension is 4.
Secondly, observe that the 4-element set $\{1,\omega_{1},\omega_{3},\omega_{1}\omega_{3}\}$ generates the $\mathbb{Q}$-vector space $\mathbb{Q}[\omega_{1},\omega_{3}]$. Thus, this set is a basis for $M_{L}$.
Lastly, define the map $\varphi:\mathbb{Z}^{4}\rightarrow M_{L}$, where $\varphi((a,b,c,d))=a\cdot 1+b\cdot\omega_{1}+c\cdot\omega_{3}+d\cdot\omega_{1}\omega_{3}$. We show that this map is an isomorphism. The injectivity of $\varphi$ is clear from the linear independence of $1$, $\omega_{1}$, $\omega_{3}$, and $\omega_{1}\omega_{3}$. The surjectivity of $\varphi$ comes from the definition of $M_{L}$. Thus, $\mathbb{Z}^{4}\cong M_{L}$. ∎
@@
Define the map $\varphi:\mathbb{Z}^{4}\rightarrow M_{L}$, where $\varphi((a,b,c,d))=a\cdot 1+b\cdot\omega_{1}+c\cdot\omega_{3}+d\cdot\omega_{1}\omega_{3}$. For a UDG $U=(V,E)$ with vertices on $M_{L}$, the matrix form of $U$ is represented as $U=[\varphi^{-1}(v_{1}),\varphi^{-1}(v_{2}),\ldots,\varphi^{-1}(v_{n})]$, where the vertex set $V=\{v_{1},v_{2},\ldots,v_{n}\}$ and the vertices are ordered.
@@
Take an arbitrary element $z=a+b\omega_{1}+c\omega_{3}+d\omega_{1}\omega_{3}\in M_{L}$ with $a,b,c,d\in\mathbb{Z}$. It is of unit length if and only if
@@
$=(a+b\omega_{1}+c\omega_{3}+d\omega_{1}\omega_{3})\overline{(a+b\omega_{1}+c\omega_{3}+d\omega_{1}\omega_{3})}$
@@
$\text{with }\mathbf{v}:=\begin{pmatrix}a\\ b\\ c\\ d\end{pmatrix}\text{ and }A:=\begin{pmatrix}1&\frac{1}{2}&\frac{5}{6}&\frac{5}{12}\\ \frac{1}{2}&1&\frac{5}{12}&\frac{5}{6}\\ \frac{5}{6}&\frac{5}{12}&1&\frac{1}{2}\\ \frac{5}{12}&\frac{5}{6}&\frac{1}{2}&1\end{pmatrix}\text{ we have }p(a,b,c,d)=\mathbf{v}^{T}A\mathbf{v}.$
@@
Remark 1.
(1) Inspired by the first preprint of this work, it was proven that in the family of lattices
$L_{k}=\{a\cdot 1+b\cdot\omega_{1}+c\cdot\omega_{k}+d\cdot\omega_{1}\omega_{k}\mid a,b,c,d\in\mathbb{Z}\},$
the maximum number of unit vectors is unbounded [7, Theorem 2.1].
(2) Let $\alpha$ be a complex root of the polynomial $p(z)=z^{4}-z^{3}-z^{2}-z+1$. Then the lattice
$L=\{a\cdot 1+b\cdot\alpha+c\cdot\alpha^{2}+d\cdot\alpha^{3}\mid a,b,c,d\in\mathbb{Z}\}$
has infinitely many unit vectors [6, Theorem 2].
We checked some of these alternative lattices, but none gave better results than the Moser lattice.
@@ 2.2. Canonization
Consider some matrix $[[p_{1}],[p_{2}],...,[p_{n}]]$ where each $[p_{i}]$ is a point on the Moser lattice, $[p_{i}]=[a_{i},b_{i},c_{i},d_{i}]=a_{i}+b_{i}\cdot\omega_{1}+c_{i}\cdot\omega_{3}+d_{i}\cdot\omega_{1}\omega_{3}$. Together, this array of points will constitute the vertex set of a planar unit distance graph. Immediately, there are $n!$ identical copies of this graph according to the permutations of this array. This may be addressed by treating this list as a set object.
@@
In the Moser lattice, we find $12$ representations of a graph through rotations by $\frac{\pi}{3}$ radians and reflections swapping the pair of generators $\omega_{1}$ and $\omega_{3}$ along with the pair $\omega_{1}\omega_{3}$ and $1$.
@@
$1=e^{0},\omega_{1}=e^{\frac{\pi i}{3}},\omega_{3}=e^{i\cdot 2\arcsin(\frac{1}{\sqrt{12}})}$
@@
$a\mapsto a\cdot\omega_{1},b\cdot\omega_{1}\mapsto-b+b\cdot\omega_{1},c\cdot\omega_{3}\mapsto c\cdot\omega_{1}\omega_{3},\text{ and }d\cdot\omega_{1}\omega_{3}\mapsto-d\cdot\omega_{3}+d\cdot\omega_{1}\omega_{3}.$
@@ 4.3.1. Operation 1: Single Edge Addition
Given a parent graph $G$ and a vertex $v\in V(G)$, generate $V(G^{\prime})=V(G)\cup v^{\prime}$ for a $v^{\prime}$ in the set of generators times ±$\pm 1$, {±$\{\pm 1,\pm\omega_{1},\pm\omega_{3},\pm\omega_{1}\omega_{3}\}$.
@@ 4.4. Parents
Finding clusters of edge-dense graphs using backtracking beam search relies also on finding subgraphs which lead to larger maximal graphs. These graphs can also therefore be found as parents of previously discovered graphs in the search. The algorithm for discovering parents is relatively simple and low-cost, and is as follows. Given a graph $G$, we obtain the parent set of $G$ by creating some $V(G^{\prime})=V(G)-v$ for all $v\in V(G)$. The set of unique canonized parents of $G$ is thus at most the size of $G$, and some subgraphs may reduce to the same canonized graph.
@@ 6.2. The Moser Ring
$(a\cdot 1+b\cdot\omega_{1}+c\cdot\omega_{3}+d\cdot\omega_{1}\omega_{3})(a^{\prime}\cdot 1+b^{\prime}\cdot\omega_{1}+c^{\prime}\cdot\omega_{3}+d^{\prime}\cdot\omega_{1}\omega_{3}),$
for all $a,b,c,d,a^{\prime},b^{\prime},c^{\prime},d^{\prime}\in\mathbb{Z}$. $M_{R}$ is an extension ring $\mathbb{Z}[\omega_{1},\omega_{3}]$, where $\omega_{1}^{2}=\omega_{1}^{2}-1$, $66\omega_{3}^{2}=10\omega_{3}-$ 6, and $\omega_{3}^{2}=\frac{5}{3}\omega_{3}-1$.
@@
$B=\{1,\omega_{1},\omega_{3},\omega_{1}\omega_{3}\},$
where $\omega_{1}=\frac{1+i\sqrt{3}}{2}$ and $\omega_{3}=\frac{5+i\sqrt{11}}{6}$. Similarly to the matrix form of UDGs with vertices in the Moser lattice, as given in Definition 2.4, every element $z\in M_{R}$ of the Moser ring can be written as $z=\frac{\sum_{b_{i}\in B}a_{i}b_{i}}{3^{k}}$ for some $(a_{0},a_{1},a_{2},a_{3},k)\in\mathbb{Z}^{5}$. Moreover, simplifying fractions, we get a unique such representation. Thus, for a UDG with $n$ vertices in the Moser ring, we could represent it by an $n\times 5$ integer matrix. All the beam search algorithms would be similar to the Moser lattice, except for the get-children function. We have a multiplication operator besides an addition operator for creating children of a UDG. Therefore, we potentially can get more children in the Moser ring than in the Moser lattice.
@@ Acknowledgements
We would like to thank Helmut Ruhland for alternative lattice ideas and investigating lattices with many unit vectors [7].
@@
There are no relevant financial or non-financial competing interests to report.